# The Design Space of Lockfiles Across Package Managers

### v1

@@ Abstract.
We perform the first comprehensive study of lockfiles across 7 popular package managers, npm, pnpm, Cargo, Poetry, Pipenv, Gradle, and Go. First, we highlight how the content and functions of lockfiles differ across package managers and ecosystems. Next, we conduct a qualitative analysis based on semi-structured interviews with 15 developers. We capture first-hand insights about the benefits that developers perceive in lockfiles, as well as the challenges they face to manage these files. Following these observations, we make 4 recommendations to further improve lockfiles, for a better developer experience.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0
1. Introduction
@@
Software reuse is an undeniable best practice in software engineering (Krueger, 1992). Software developers reuse third-party libraries for commodity features such as logging, cryptography, arithmetic, or I/O. Reuse helps developers save time, pool quality assurance efforts, and focus on the core, application specific features of their project (Mohagheghi and Conradi, 2007). Yet, software reuse at scale introduces major engineering challenges (Cox, 2019). The use of third-party libraries, including the transitive ones, often results in large dependency trees. Managing these increasingly complex dependency trees is hard because developers have to resolve compatible versions across the whole tree (Rausch et al., 2017), stay on top of updates to patch bugs and vulnerabilities (Ladisa et al., 2023), and ensure that new changes do not break the build (Frank et al., 2024).
@@
Package managers are tools that assist developers in the management of reused libraries (Spinellis, 2012). They automate and simplify the resolution, download, configuration, installation, and update of dependencies. When using a package manager, developers define the dependencies of a project in a dependency specification file. Then, the package manager determines compatible versions of these dependencies and all indirect dependencies, fetches them from a package registry, and installs them to further build the project. During this process, some package managers save the exact list of resolved dependencies and their versions into a special file: the lockfile (Yu et al., 2024).
@@
Lockfiles are widely used throughout the software development lifecycle, supporting reproducible builds (Lamb and Zacchiroli, 2022a), defending the software supply chain (Liu et al., 2022), and contributing to improved build performance (Aïdasso et al., 2025). However, lockfiles also introduce additional complexity, and their maintenance requirements are not fully understood, a subject hardly touched by software engineering research. Lockfiles have been mentioned as security best practices (Kabir et al., 2022), important for software supply chain security (Williams et al., 2025), related to Software Bill of Materials (SBOM) generation (Bi et al., 2024), and notable for new dependency management tools (Cleare and Iacob, 2018). However, none of these works provide an in-depth analysis of lockfiles, why they exist and what they should contain. To our knowledge, there is no systematic analysis and comparison of the different strategies in the field to generate and evolve lockfiles. There is no grounded evidence on how developers handle and make the best out of their lockfiles.
@@
In this study, we address this critical gap in the software engineering literature by conducting a comprehensive analysis of lockfiles across 7 package managers. These package managers serve the top programming languages of 2024 (Cass, 2024): npm and pnpm for Javascript, Cargo for Rust, Poetry and Pipenv for Python, Gradle for Java, and Go for Golang. We analyze the official documentation and implementation of these package managers, and compare the strategies they implement to generate and maintain lockfiles. We then quantitatively evaluate lockfile usage in the wild, by collecting GitHub projects and their lockfile commit rates. To complement this analysis, we conduct interviews with 15 developers with diverse backgrounds and experiences. They contribute to active open source projects spanning various domains such as networking, gaming, CLI, and APIs, and use the package managers under study. Through these interviews, we gather valuable first-hand insights into the developers’ motivations for using or avoiding lockfiles, as well as the challenges they face when working with them.
@@
Our paper is the first study that highlights how the functionalities and usage of lockfiles differ across package managers and ecosystems, resulting in varying levels of developer satisfaction. While some approaches, such as the Go strategy for lockfiles, encourage strong developer engagement, others, such as Gradle’s approach, fail to address key software engineering challenges. Our findings clearly show that a simple, principled approach to the lockfile lifecycle leads to a better developer experience. We also gather evidence of the challenges that developers face with lockfiles, such as limited human readability, difficulty locking indirect dependencies, and delays in receiving dependency updates. Overall, developers generally trust lockfiles and view them as a valuable tool for improving the efficiency and security of dependency management. We conclude our study with 4 recommendations for developers and package manager maintainers to encourage the adoption of lockfiles in all projects.
All data and interview protocols used in this study are publicly available: https://github.com/chains-project/lockfiles-comprehensive-study. In summary, our key contributions are,
@@
  • a comprehensive comparison of lockfiles across 7 top package managers, with a deep analysis of the key dimensions of the lockfile problem space.
@@
  • a regimen of 4 recommendations to further improve lockfiles for dependency management, to inform future research in the field as well as engineering of the future generation of package managers.
2. Definitions and Context
@@
2.1. Definitions
Figure 1. A canonical workflow of interactions between a developer, the dependency specification file, the package manager, the lockfile, and the package registry.
@@ 2.1. Definitions
Table 1. The 7 package managers analyzed in this study, including their associated programming languages, registries, manifests, and lockfiles.
@@
Gradle Java central.sonatype.com build.gradle gradle.lockfile
@@
2.2. Considered Package Managers
We select 7 popular package managers for our study based on the popularity of programming languages in 2024 (Cass, 2024). From the ten most popular languages, we identify those with widely used package managers that support lockfiles: JavaScript, TypeScript, Python, Java, and Go. We add Rust due to its significant upward trend. Table 1 outlines the 7 package managers analyzed in this study, along with their associated language, registries, dependency specification files, and lockfiles.
@@ 2.2. Considered Package Managers
Table 2. Details about the exact versions of the 7 package managers under study.
Package manager Analyzed Version Lines of code Released versions
@@
Listing 1: An excerpt from the dependency specification file (package.json) of the project asyncapi/modelina
@@
9 "integrity": "sha512-WZyqzPLrWctQB8qB8uGivvzWgvtcVqlmR...",
@@
2buf.build/gen/go/bufbuild/bufplugin/protocolbuffers/go v1.36.3-.../go.mod h1:MYDFm9IHRP085R5Bis68mLc0mIqp5Q27Uk...
@@
3. Study of Lockfile Differences
3.1. Research Questions
  1. RQ1:
    How do package managers generate, update, and structure lockfiles?
  2. RQ2:
    In this RQ, we quantitatively evaluate how developers of different ecosystems include lockfiles in version control. Including lockfiles in version control systems is important to achieve all benefits lockfiles have to offer, including deterministic builds (Goswami et al., 2020). We base our evaluation on lockfile commit data collected from open source GitHub projects created within the past 5 years (from 2019 to 2024). This quantitative analysis helps consolidate our observations in RQ1 about the usability and usefulness of lockfiles.
3.2. Methodology for RQ1
@@
Source code analysis: To understand the workflows of lockfile-related tasks for each package manager, we analyze their implementation within their source code repository. We begin by locating the main function that is invoked after executing an install or build command. We then analyze the functions called by this main function, manually tracing their execution flow. We further search the project using regular expressions for terms such as lockfile, locked, frozen, and dependency resolution. The sections of source code, analyzed for each package manager are summarized in Table 3. For instance, in the case of npm, we study the repository npm/cli, focusing on the source code for install and ci actions, dependency tree resolution, and lockfile validation.
@@
Experiments with sample projects: To observe the dependency resolution behavior, with and without lockfiles, under different command-line options, we conduct experiments using sample projects for each package manager. First, we study the generation and validation of lockfiles under default configurations using the install or build command. Next, we experiment with various flags, such as --deploy or --locked, which trigger different lockfile validation behaviors. If we consider Gradle, as shown in Table 3, we build the project both with and without the --write-locks option and also with and without defining lock states in the dependency specification file. Further, we experiment with different locking modes strict, lenient, and default.
Table 3. Lockfile features for the 7 package managers under study. We study their source code, documentation, and experiment with sample projects.
Package manager Sections of source code Sections of documentation Commands run
npm CLI install and ci actions, dependency tree resolution (within arborist module), lockfile validation CLI commands npm install, npm ci
pnpm lockfile module, install action CLI commands - Manage dependencies pnpm i, pnpm i --frozen-lockfile, pnpm i --offline
Cargo install action, lockfile functions Build commands, manifest commands, Cargo.toml vs Cargo.lock cargo build, cargo build --locked, cargo build --offline, cargo build --frozen
Poetry installer class, locker class Installing dependencies poetry install, poetry sync
Pipenv command.py, install.py, locking.py, lock.py Pipenv commands, Pipfile & Pipfile.lock pipenv install, pipenv install --deploy, pipenv sync
Gradle locking package, dependencies package Locking versions gradle build, gradle build --write-locks (both with and without lock states in different modes; strict, lenient, and default)
Go modcmd, modload, modfetch, modget, mvs Managing dependencies Go Modules Reference go.mod file reference go mod init, go mod tidy, go get, go mod edit, go build
3.3. Methodology for RQ2
@@
Excluding projects with no dependencies: For projects that use a package manager but do not declare any dependencies, lockfiles do not provide any added value. Therefore, we collect the number of direct dependencies specified in each dependency specification file and evaluate lockfile usage only for projects that have at least one declared direct dependency. In multi-module projects, if the main dependency specification file in the root folder does not specify any dependencies, we check the dependency specification files of sub-modules. If any sub-module declares third-party dependencies, we keep the project.
3.4. Results for RQ1: How do package managers generate, update, and structure lockfiles?
@@ 3.4.1. Lockfile Content
Link to Source NPM and Cargo include a provenance link to source code. This is excellent for future rebuilding efforts (Lamb and Zacchiroli, 2022b). In terms of naming, npm CLI uses the field resolved for source links, while Cargo uses source. In Go, the source can be inferred from the module name declaration itself, as we see from the link to the GitHub repository for the module text in 11. Gradle, pnpm, Poetry, and Pipenv do not include a direct link to the source of the downloaded packages.
@@
Version ranges Cargo omits the version ranges of indirect dependencies (when listed under each parent dependency) unless multiple versions of the same dependency are required. When a version is included, Cargo specifies the exact resolved version. In contrast, npm, pnpm, and Poetry include version ranges for indirect dependencies under a parent dependency, reflecting the same constraints set in the dependency specification files. A better practice would be to display the exact resolved version.
Additional metadata Some package managers include additional details such as language requirements (e.g. Poetry, Pipenv), operating systems (e.g. npm, Cargo), or kernel compatibility (e.g. npm). Moreover, npm CLI includes in the lockfile extra metadata such as license details and funding information, directly copied from package.json (5). Although recording as much information as possible for each dependency may seem a good choice, it also makes the lockfiles lengthier and harder to be reviewed by human developers.
@@
3.4.2. Lockfile Lifecycle
Table 4. Lockfile life cycles across package managers. All behaviors shown here are observed when invoking projects build under the default configuration. Special flags required to achieve the corresponding behavior are given within brackets.
Package manager Lockfile generation Lockfile based dependency resolution Lockfile enforcement - specification file mismatches
npm CLI by default by default optional (must use npm ci)
pnpm by default by default optional (only with --frozen-lockfile)
Cargo by default optional (only with --locked or --frozen flags) optional (only with --locked or --frozen flags)
@@ 3.4.2. Lockfile Lifecycle
Pipenv by default optional (can use -deploy flag or sync action) optional (can use -deploy flag or sync action)
Gradle optional (only with –write-locks flag and specified lock states) by default by default
@@
1"node_modules/@eslint/eslintrc": { 2 "version": "2.1.4", 3 "resolved":"https://registry.npmjs", 4 "integrity":"sha512-269Z39MS6wVJts...", 5 "dev": true, 6 "dependencies": { "ajv": "^6.12.4", }, 7 "engines": { 8 "node": "^12.22.0" 9 }, 10 "funding": { 11 "url": "" 12 } 13} Listing 5: package-lock.json of the project asyncapi/modelina
1@vitejs/plugin-vue@5.0.4: 2 resolution:integrity:sha512-WS3hevEsz... 3 engines: {node: ^18.0.0 $|$ >=20.0.0} 4 peerDependencies: vite: ^5.0.0 Listing 6: pnpm-lock.yaml of the project csfive/composing-programs-zh
1[[package]] 2name = "Inflector" 3version = "0.11.4" 4 source = "registry+https://github.com..." 5checksum="fe438c63458706e0..." 6 dependencies = [ "lazy_static" ] Listing 7: Cargo.lock of the project leptos-rs/cargo-leptos
1[[package]] 2name = "jinja2" 3version = "3.1.4" 4description = "" 5optional = false 6python-versions =">=3.7" 7files = 8 [{file ="jinja.tar.gz", 9 hash="sha:2564a3aee7..."}] 10 [package.dependencies] MarkupSafe = ">=2.0" 11[package.extras] 12i18n = ["Babel (>=2.7)"] Listing 8: poetry.lock of the project mingrammer/diagrams
1"plexapi": { 2"hashes":["sha=2565af051..."], 3"index": "pypi", 4"markers": "python_version>=’3.9’", 5"version": "==4.16.1" 6}, Listing 9: Pipfile.lock of the project Drazzilb08/daps
1require ( 2 github.com/kr/text v0.2.0 //indirect 3) 4-------------------------------------- 5github.com/kr/text v0.2.0 h1:5Nx0Ya0ZqY2ygV366QzturHI13Jq95ApcVaJBhp 6github.com/kr/text v0.2.0/go.mod h1:eLer722TekiGuMkidMxCpM04lWEeraHUUmBw8l2 Listing 11: go.mod and go.sum (below the dashed line) of the project piplabs/story
Figure 2. Example dependency representations extracted from lockfiles (or a mod file) across 7 package managers. Dependency versions are highlighted in yellow, source links in blue, checksums in green, and indirect dependencies in pink. Some details are omitted for brevity and clarity.
@@
Lockfile enforcement. Lockfile enforcement is the property of failing the build if the lockfile conflicts with the dependency specification file. Here, a conflict means a discrepancy in the checksum or the resolved version. In the presence of such a conflict, npm CLI and pnpm resolve versions compatible with the dependency specification file, overriding the lockfile silently (no enforcement). In contrast, some package managers fail the build in case of a conflict, thus enforcing the lockfile. This is the case for Poetry and Gradle where the lockfile is enforced by default, and an error is raised if conflicts are found. In npm, only npm ci command enforces the lockfile, and in pnpm, enforcement only happens with the --frozen-lockfile flag. In Cargo and Pipenv, the same flags required for lockfile based dependency resolution (--locked and --deploy) also achieve lockfile enforcement. For Go, lockfile enforcement happens by default because it does not have a separate developer maintained dependency specification file, and the checksums in go.sum are automatically validated.
@@
Findings from RQ1: Lockfiles address a fundamental problem of software engineering: knowing and verifying the exact versions of components required to build a software system. Cargo and Go lockfiles are feature complete. npm CLI lockfiles tend to be unnecessarily verbose due to excessive metadata. Gradle lockfiles are overly simplistic as they skip mandatory details such as checksums. Simple, streamlined lockfile processes such as Go’s are fool proof and recommended as blueprint for the future. Our findings are useful for developers to learn about lockfiles and for package manager maintainers to correctly design and implement effective lockfile features.
Table 5. The most common types of information included in lockfiles
Package manager Resolved package versions Package checksums Package source Direct/indirect differentiation Additional metadata
npm CLI $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (under each dependency), Dependency scopes, OS, engine, cpu, kernel requirements, Licenses, Funding information
@@
Cargo $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (under each dependency), Language, OS, engine requirements
@@
3.5. Results for RQ2: To what extent do open-source projects include lockfiles in version control?
Table 6. GitHub projects analyzed to assess the presence of lockfiles in the version control system. The third column shows projects that started including lockfiles in version control within six months of project creation. The final column shows the number of projects that currently commit lockfiles.

### v2

@@ Abstract
We perform the first comprehensive study of lockfiles across 7 popular package managers, npm, pnpm, Cargo, Poetry, Pipenv, Gradle, and Go. First, we highlight the wide variety of design decisions that package managers make, regarding the generation process as well as the content of lockfiles. Next, we conduct a qualitative analysis based on semi-structured interviews with 15 developers. We capture first-hand insights about the benefits that developers perceive in lockfiles, as well as the challenges they face to manage these files. Following these observations, we make 5 recommendations to further improve lockfiles, for a better developer experience.
1 Introduction
@@
Software reuse is an undeniable best practice in software engineering [1]. Software developers reuse third-party libraries for commodity features such as logging, cryptography, arithmetic, or I/O. Reuse helps developers save time, pool quality assurance efforts, and focus on the core, application specific features of their project [2]. Yet, software reuse at scale introduces major engineering challenges [3]. The use of third-party libraries, including the transitive ones, often results in large dependency trees. Managing these increasingly complex dependency trees is hard because developers have to resolve compatible versions across the whole tree [4], stay on top of updates to patch bugs and vulnerabilities [5], and ensure that new changes do not break the build [6].
@@
Package managers are tools that assist developers in the management of reused libraries [7]. They automate and simplify the configuration, resolution, installation, and update of dependencies. When using a package manager, developers define the dependencies of a project in a dependency specification file. Then, the package manager determines compatible versions of these dependencies and all indirect dependencies, fetches them from a package registry, and installs them to further build the project. During this process, some package managers save the exact list of resolved dependencies and their versions into a special file: the lockfile [8].
@@
Lockfiles are widely used throughout the software development lifecycle, supporting reproducible builds [9], defending the software supply chain [10], and contributing to improved build performance [11]. However, lockfiles also introduce additional complexity, and their maintenance requirements are not fully understood, a subject hardly touched by software engineering research. Lockfiles have been mentioned as security best practices [12], important for software supply chain security [13], related to Software Bill of Materials (SBOM) generation [14, 8], and notable for new dependency management tools [15]. However, none of these works provide an in-depth analysis of lockfiles, why they exist and what they should contain. To our knowledge, there is no systematic analysis and comparison of the different strategies in the field to generate and evolve lockfiles. There is no grounded evidence on how developers perceive the benefits and challenges of maintaining lockfiles.
@@
In this study, we address this critical gap in the software engineering literature by conducting a comprehensive analysis of lockfiles across 7 package managers. These package managers serve the top programming languages of 2024 [16]: npm and pnpm for Javascript, Cargo for Rust, Poetry and Pipenv for Python, Gradle for Java, and Go for Golang. We analyze the implementation of these package managers, to unveil the large design space for the generation, maintenance and content of lockfiles. We then quantitatively evaluate lockfile usage in the wild, by collecting GitHub projects and their lockfile commit rates. To complement this analysis, we conduct interviews with 15 developers with diverse backgrounds and experiences. They contribute to active open source projects spanning various domains such as networking, gaming, CLI, and APIs, and use the package managers under study. Through these interviews, we gather valuable and grounded first-hand insights into the developers’ motivations for using or avoiding lockfiles, as well as the challenges they face when working with them.
@@
Our paper is the first study that highlights how the functionalities and usage of lockfiles differ across package managers and ecosystems, resulting in varying levels of developer satisfaction. While some approaches, such as the Go strategy for lockfiles, encourage strong developer engagement, others, such as Gradle’s approach, fail to address key software engineering challenges. Our findings clearly show that a simple, principled approach to the lockfile lifecycle leads to a better developer experience. We also gather evidence of the challenges that developers face with lockfiles, such as limited human readability, difficulty locking indirect dependencies, and delays in receiving dependency updates. Overall, developers generally trust lockfiles and view them as a valuable tool for improving the efficiency and security of dependency management. We conclude our study with 5 recommendations to encourage the adoption of lockfiles in all projects. Our findings are directly actionable by package manager users and by package manager maintainers. Our results assist package manager users in understanding lockfiles and choosing a package manager that better suits their project requirements. Our findings help package manager maintainers improve their lockfile feature, which in turn can positively impact a large community of developers. All data and interview protocols used in this study are publicly available: https://github.com/chains-project/lockfiles-comprehensive-study. In summary, our key contributions are,
@@
  • the first comprehensive comparison of lockfiles across the 7 top package managers, with a deep analysis of the key dimensions of the lockfile problem space: the lockfile lifecycle, and content of the generated lockfile
@@
  • a regimen of 2 recommendations for package manager users and 3 recommendations for package manager developers to further improve lockfile usage for dependency management, to inform future research in the field as well as engineering of the future generation of package managers.
2 Definitions and Context
@@
2.1 Definitions
Figure 1: A canonical workflow of interactions between a developer, the dependency specification file, the package manager, the lockfile, and the package registry.
@@ 2.1 Definitions
Table 1: The 7 package managers analyzed in this study, including their associated programming languages, registries, dependency specifications, and lockfiles.
@@
Gradle Java central.sonatype build.gradle gradle.lockfile
@@
2.2 Considered Package Managers
We select 7 popular package managers for our study based on the popularity of programming languages in 2024 [16]. From the ten most popular languages, we identify those with widely used package managers that support lockfiles: JavaScript, TypeScript, Python, Java, and Go. We add Rust due to its significant upward trend. For each language, we select the most commonly used package manager with lockfile support: npm for JavaScript/TypeScript, Poetry for Python, Cargo for Rust, Gradle for Java, and Go’s built-in toolchain. When a language has multiple package managers, we add a second one: pnpm for JavaScript and Pipenv for Python. Table 1 outlines the 7 package managers analyzed in this study, along with their associated languages, registries, dependency specification files, and lockfiles.
@@ 2.2 Considered Package Managers
Table 2: Details about the exact versions of the 7 package managers under study.
Package manager Analyzed Version Lines of code Versions
@@
3 Comprehensive Comparison of Lockfile Design Decisions
3.1 Research Questions
  1. RQ1:
    How do package managers generate, update, and structure lockfiles?
  2. RQ2:
    In this RQ, we quantitatively evaluate how developers of different ecosystems include lockfiles in version control. Including lockfiles in version control systems is important to achieve all benefits lockfiles have to offer, including deterministic builds [17]. We base our evaluation on lockfile commit data collected from open source GitHub projects created within the past 5 years (from 2019 to 2024). This quantitative analysis helps consolidate our observations in RQ1 about the usability and usefulness of lockfiles.
@@ 2. RQ2:
Listing 1: An excerpt from the specification file (package.json) of the project asyncapi/modelina
@@
9 "integrity": "sha512-WZyqzPLrWctQB8qB8uGivvzWg...",
@@
2buf.build/gen/go/bufbuild/bufplugin/protocolbuffers/go v1.36.3-.../go.mod h1:MYDFm9IHRP085R5Bis68mLc0mI...
@@
3.2 Methodology for RQ1
@@
Source code analysis: To understand the workflows of lockfile-related tasks for each package manager, we analyze their implementation within their source code repository. We begin by locating the main function that is invoked after executing an install or build command. We then analyze the functions called by this main function, manually tracing their execution flow. We further search the project using regular expressions for the terms: lockfile, locked, frozen, and dependency resolution. The sections of source code, analyzed for each package manager are summarized in Table 3. For instance, in the case of npm, we study the repository npm/cli, focusing on the source code for install and ci actions, dependency tree resolution, and lockfile validation.
@@
Table 3: Lockfile features for the 7 package managers under study. We study their source code and documentations.
Package manager Sections of source code Sections of documentation
npm CLI install and ci actions, dependency tree resolution (within arborist module), lockfile validation CLI commands
pnpm lockfile module, install action CLI commands - Manage dependencies
Cargo install action, lockfile functions Build commands, manifest commands, Cargo.toml vs Cargo.lock
Poetry installer class, locker class Installing dependencies
Pipenv command.py, install.py, locking.py, lock.py Pipenv commands, Pipfile & Pipfile.lock
Gradle locking package, dependencies package Locking versions
Go modcmd, modload, modfetch, modget, mvs Managing dependencies Go Modules Reference go.mod file reference
3.3 Methodology for RQ2
@@
Excluding projects with no dependencies: For projects that use a package manager but do not declare any dependencies, lockfiles do not provide any added value. Therefore, we collect the number of dependencies specified in each dependency specification file and evaluate lockfile usage only for projects that have at least one declared dependency. In multi-module projects, if the main dependency specification file in the root folder does not specify any dependencies, we check the dependency specification files of sub-modules. If any sub-module declares third-party dependencies, we keep the project.
3.4 Results for RQ1: How do package managers generate, update, and structure lockfiles?
@@ 3.4.1 Lockfile Content
Link to Source NPM and Cargo include a provenance link to source code. This is excellent for future rebuilding efforts [9]. In terms of naming, npm CLI uses the field resolved for source links, while Cargo uses source. In Go, the source can be inferred from the module name declaration itself, as we see from the link to the GitHub repository for the module text in 11. Gradle, pnpm, Poetry, and Pipenv do not include a direct link to the source of the downloaded packages.
@@
Version ranges Cargo omits the version ranges of indirect dependencies (when listed under each parent dependency) unless multiple versions of the same dependency are required. When a version is included, Cargo specifies the exact resolved version. In contrast, npm, pnpm, and Poetry include version ranges for indirect dependencies under a parent dependency, reflecting the same constraints set in the dependency specification files. In addition to version ranges, it is valuable for developers to have straightforward access to the exact resolved versions.
Additional metadata Some package managers include additional details such as language requirements (e.g. Poetry, npm), operating systems (e.g. npm, Cargo) or dependency scope (e.g. npm, Gradle). Moreover, npm CLI includes in the lockfile extra metadata such as license details and funding information, directly copied from package.json (5). Although recording as much information as possible for each dependency may seem a good choice, it also makes the lockfiles lengthier and harder to be reviewed by human developers.
@@
1"node_modules/@eslint/eslintrc": { 2"version": "2.1.4", 3"resolved":"https
:registry.npmjs"
4"integrity":"sha512-269Zhnju8...", 5"dev": true, 6 "dependencies": { "ajv": "^6.12.4", }, 7"engines": { 8 "node": "^12.22.0" 9 }, 10"funding": { 11 "url": "" 12 } 13}
Listing 5: package-lock.json of the project asyncapi/modelina
1@vitejs/plugin-vue@5.0.4: 2resolution:integrity:sha512-WS3R45... 3engines: {node: ^18.0.0 $|$ >=20.0.0} 4 peerDependencies: vite: ^5.0.0 Listing 6: pnpm-lock.yaml of the project csfive/composing-programs-zh
1[[package]] 2name = "Inflector" 3version = "0.11.4" 4 source = "registry+https://..." 5checksum="fe438t67y6..." 6 dependencies = [ "lazy_static" ] Listing 7: Cargo.lock of the project leptos-rs/cargo-leptos
1[[package]] 2name = "jinja2" 3version = "3.1.4" 4description = "" 5optional = false 6python-versions =">=3.7" 7files = 8[{file =".tar.gz", 9hash="sha:25646778..."}] 10 [package.dependencies] MarkupSafe = ">=2.0" Listing 8: poetry.lock of the project mingrammer/diagrams
1"plexapi": { 2"hashes":["sha=256..."], 3"index": "pypi", 4"markers": "python_version>=’3.9’", 5"version": "==4.16.1" 6}, Listing 9: Pipfile.lock of the project Drazzilb08/daps
1require ( 2 github.com/kr/text v0.2.0 //indirect 3) 4----------------------------- 5github.com/kr/text v0.2.0 h1:5Nx0Ya0ZqY2ygV366Qz... 6github.com/kr/text v0.2.0/go.mod h1:eLer722TekiGuMkidMxCpM0... Listing 11: go.mod and go.sum (below the dashed line) of the project piplabs/story
Figure 2: Example dependency representations extracted from lockfiles (or a mod file) across 7 package managers. Dependency versions are highlighted in yellow, source links in blue, checksums in green, and indirect dependencies in pink. Some details are omitted for brevity and clarity.
3.4.2 Lockfile Lifecycle
Table 4: Lockfile life cycles across package managers. All behaviors shown here are observed when invoking projects build under the default configuration. Special flags required to achieve the corresponding behavior are given within brackets.
Package manager Lockfile generation Lockfile-based resolution Lockfile enforcement on spec file mismatch
npm CLI by default by default optional (npm ci)
pnpm by default by default optional (--frozen-lockfile)
Cargo by default optional (--locked, --frozen) optional (--locked, --frozen)
@@ 3.4.2 Lockfile Lifecycle
Pipenv by default optional (--deploy, sync) optional (--deploy, sync)
Gradle optional (--write-locks with specified lock states) by default by default
@@
Lockfile enforcement. Lockfile enforcement is the property of failing the build if the lockfile conflicts with the dependency specification file. Here, a conflict means a discrepancy in the checksum or the resolved version. In the presence of such a conflict, npm CLI and pnpm resolve versions compatible with the dependency specification file, overriding the lockfile silently (no enforcement). In contrast, some package managers fail the build in case of a conflict, thus enforcing the lockfile. This is the case for Poetry and Gradle where the lockfile is enforced by default, and an error is raised if conflicts are found. In npm, only npm ci command enforces the lockfile, and in pnpm, enforcement only happens with the --frozen-lockfile flag. In Cargo and Pipenv, the same flags required for lockfile based dependency resolution (--locked and --deploy) also achieve lockfile enforcement. For Go, lockfile enforcement happens by default because it does not have a separate developer maintained dependency specification file, and the checksums in go.sum are automatically validated when a package gets downloaded.
@@
Findings from RQ1: Lockfiles address a fundamental problem of software engineering: knowing and verifying the exact versions of components required to build a software system. Cargo, Go, and npm CLI lockfiles contain all essential dependency information: resolved package versions, package checksums, package sources. However, npm CLI lockfiles are unnecessarily verbose due to excessive metadata. Gradle lockfiles are overly simplistic as they skip mandatory information such as checksums. Simple, streamlined lockfile processes such as Go’s are fool proof and recommended as blueprint for the future.
Table 5: The most common types of information included in lockfiles
Package manager Resolved package versions Package checksums Package source Direct/ indirect difference Additional metadata
npm CLI $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (for each dependency), Dependency types, OS, engine, cpu, requirements, Licenses, Funding information
@@
Cargo $usym1F5F8 \usym$$\usym1F5F8$ Indirect dependencies (for each dependency), Language, OS, engine requirements
@@
3.5 Results for RQ2: To what extent do open-source projects include lockfiles in version control?
Table 6: GitHub projects analyzed to assess the presence of lockfiles in the version control system. The third column shows projects that started including lockfiles in version control within six months of project creation. The final column shows the number of projects that currently commit lockfiles.